OpenStudy (anonymous):

looking for help finding factors that when multipied equal 430 but when added equal 76. Can anyone help?

OpenStudy (anonymous):

Are you sure it isn't 420? If it is, it is 6 and 70

OpenStudy (anonymous):

no unfortunely it is 430.

OpenStudy (bahrom7893):

x*y = 430 x+y = 76

OpenStudy (bahrom7893):

x*y = 430 y = 76-x

OpenStudy (bahrom7893):

x*(76-x)=430 y = 76-x

OpenStudy (bahrom7893):

what do u mean factors? do u mean just 2 numbers?

OpenStudy (bahrom7893):

76x-x^2 = 430 y = 76-x

OpenStudy (anonymous):

like a . b = 430 and then a+b=76

OpenStudy (bahrom7893):

OMG this place crashes too much so: -x^2 +76x - 430 = 0

OpenStudy (bahrom7893):

let's see if discriminant works: D = 76^2 - 4 * (-1) * (-430) = 5776 - 1720 = 4056

OpenStudy (bahrom7893):

Uhmm.. r us sure the problem is right? Because radical D is 63.68673331, etc..

OpenStudy (bahrom7893):

i'm still thinking that blexting was right.. the question is supposed to have a 420.. Nobody in their right mind would assign a problem with decimals.. well unless ur dealin with physics

OpenStudy (anonymous):

ok maybe i went about this the wrong way. i am look for an answer for -2x squared+76x+ 430. supposed to factor by grouping i think. trying to figure it out might have messed u

OpenStudy (anonymous):

oops mean up*

OpenStudy (bahrom7893):

I think this way it makes sense.. and if u can't factor use quadratics. it's much easier

OpenStudy (bahrom7893):

-2x^2 + 76x + 430 = 0 by the way is it + 430 or - 430?

OpenStudy (bahrom7893):

D = 76^2 - 4 * (-2) * 430 = 5776 + 3440 = 9216

OpenStudy (bahrom7893):

x1 = (-76-sqrt(9216))/(2*(-2)) = (-76 - 96)/-4 = 43

OpenStudy (bahrom7893):

x2 = (-76+sqrt(9216))/(2*(-2)) = (-76 + 96)/-4 = 20/-4 = -5

OpenStudy (bahrom7893):

So to factor ur equation you would have: (x-43)(x+5)=0

OpenStudy (bahrom7893):

Please click on become a fan if I helped, I really want to get to the next level!! Thanks =)

OpenStudy (anonymous):

First of all if -2x^2 + 76x + 430 = 0 is the equation Then the first thing you do is to factor out the -2 (common factor) -2(x^2 - 38x - 215) = 0 Then you would be looking for factors of 215 that subtract to get 38 Which would be 43 and 5 So it would factor into -2(x - 43)(x + 5) = 0